# DeCoF: Generated Video Detection via Frame Consistency

### v2

@@ A.2 Dataset Evaluation
As part of this research, we present a comprehensive analysis of the GVF dataset, which includes an analysis of the attribute statistics of prompts, as well as an assessment of the different generation models in terms of the spatial content, spatial attributes, temporal content, and temporal attributes according to the FETV (Liu et al., 2023).
@@
Attribute statistics of prompts. From the four aspects mentioned before, the main content of space and time and the main attributes of space and time, we provide attribute statistics of prompts of the data set. As the figure 10 shows, the various attributes of prompts are different because one prompt may have multiple attributes at the same time. At the same time, different attributes have different difficulties in generating videos, and the harm they cause in a trust crisis is also different. You may ask why time main content and time attributes are so rare. In fact, not every prompt in the MSVD and MSR-VTT datasets has a time description, and we only ensure the existence of this type of prompt in the data set. Here, we also provide some specific examples of prompts, As shown in the table 4.
Evaluation of different generation models. Here, we present our assessment of four generation models according to the FETV standard (Liu et al., 2023). We assess the generated videos based on four perspectives: (i) Static quality focuses on the visual quality of single video frames. (ii) Temporal quality focuses on the temporal coherence of video frames. (iii) Overall alignment measures the overall alignment between a video and the given text prompt. (iv) Fine-grained alignment measures the video-text alignment regarding specific attributes.
@@
Appendix B Parameter Description
@@
Training details. During the training process, we only train on one sub-dataset of the GVF dataset and test on the other sub-datasets. The numbers of the training set, validation set, and test set are 771, 97, and 96, respectively. We freeze the parameters of CLIP:ViT (Dosovitskiy et al., 2020), making it a pure mapping function, and only train our model using the SGD optimizer with an initial learning rate of 0.001 and a momentum coefficient of 0.9.
Data pre-processing. All the following experiments are performed on our GVF dataset. In order to better detect the videos generated by the unseen generate model, we only train the model on the sub-dataset text2video-zero and test on the remaining three sub-datasets Show1 (Zhang et al., 2023), ModelScope (Wang et al., 2023a) and ZeroScope (Sterling., June 2023) Since the resolution and length of the videos generated by the four generative models are inconsistent, We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224. During the training process, we first resized the image to 256×256 according to bilinear interpolation and then randomly cropped with the size of 224 × 224. Some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
@@
Mapping Function Pretrain Tasks Tested generate video models Total
@@
Text2Video-zero Show1 ModelScope ZeroScopre
@@
ViT-B/16 ImageNet
Resnet50 ImageNet
Appendix C Different Mapping Functions
First, we discussed the two different tasks of ImageNet and CLIP. Interestingly, the encoder trained on ImageNet caused this problem not to converge. We considered two reasons: one is the size of the training data, and the other is Due to the difference in the task itself; we do not want the model to classify images into specific items or biological categories but rather cluster similar images. At the same time, we compared the impact of different model frameworks. Due to issues such as the number of model parameters, Resnet trained on the CLIP task is not suitable as a mapping function. Overall, the mapping function largely determines the effectiveness of our approach.
Appendix D More Visual Examples

### v3

@@ B.2 Dataset Evaluation
As part of this research, we present a comprehensive analysis of the GVF dataset, which includes an analysis of the attribute statistics of prompts, as well as an assessment of the different generation models in terms of the spatial content, spatial attributes, temporal content, and temporal attributes according to the FETV [25].
@@
Attribute statistics of prompts. From the four aspects mentioned before, the main content of space and time and the main attributes of space and time, we provide attribute statistics of prompts of the data set. As the figure 10 shows, the various attributes of prompts are different because one prompt may have multiple attributes at the same time. At the same time, different attributes have different difficulties in generating videos, and the harm they cause in a trust crisis is also different. Here, we also provide some specific examples of prompts, As shown in the table 4.
@@
Evaluation of different generation models. Here, we present our assessment of four generation models according to the FETV standard [25]. We assess the generated videos based on four perspectives: (i) Static quality focuses on the visual quality of single video frames. (ii) Temporal quality focuses on the temporal coherence of video frames. (iii) Overall alignment measures the overall alignment between a video and the given text prompt. (iv) Fine-grained alignment measures the video-text alignment regarding specific attributes.
@@
Appendix C Parameter Description
@@
Training details. During the training process, we only train on one sub-dataset of the GVF dataset and test on the other sub-datasets. The numbers of the training set, validation set, and test set are 771, 96, and 97, respectively. We freeze the parameters of CLIP:ViT [8], making it a pure mapping function, and only train our model using the SGD optimizer with an initial learning rate of 0.001 and a momentum coefficient of 0.9. All experiments were completed on two RTX 3090 graphics cards.
Data pre-processing. All the following experiments are performed on our GVF dataset. In order to better detect the videos generated by the unseen generate model, we only train the model on the sub-dataset text2video-zero and test on the remaining three sub-datasets Show1 [51], ModelScope [46] and ZeroScope [42] Since the resolution and length of the videos generated by the four generation models are inconsistent, We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224. During the training process, we first resized the image to 256×256 according to bilinear interpolation and then randomly cropped with the size of 224 × 224. Some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
@@
Appendix D Different Mapping Functions
First, we discussed the two different tasks of ImageNet and CLIP. Compared to ImageNet’s image classification task, CLIP’s image encoder trained on text image task alignment is a better mapping function. We considered two reasons: one is the size of the training data, and the other is due to the difference in the task itself; We do not want the model to classify images into specific items or biological categories but rather cluster similar images. At the same time, we compared the impact of different model frameworks: Due to issues such as the number of model parameters, Resnet trained on the CLIP task is not suitable as a mapping function.
Overall, the mapping function largely determines the effectiveness of our approach.
@@ Appendix D Different Mapping Functions
Mapping Function Pretrain Tasks Video generation models Total
@@
Text2Video-zero Show1 ModelScope ZeroScope
@@
ViT-B/16 ImageNet 98.00 99.79 65.39 87.65 54.70 76.39 55.57 76.1 68.41 84.98
Resnet50 ImageNet 99.5 99.96 54.2 81.9 51.43 67.97 49.93 58.65 63.76 77.12
Appendix E The Impact of Video Frame Number
Here we discuss the impact of the number of video frames used for detection.
Table 6: A comparison of the impact of different frame numbers.
Training dataset Frame Numbers Video generation models Total
Avg.
Text2Video-zero Show1 ModelScope ZeroScope
ACC AP ACC AP ACC AP ACC AP ACC AP
Text2Video-zero 8 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
Show1 8 98.50 99.77 97.00 99.33 90.25 97.96 93.75 99.08 94.88 99.04
ModelScope 8 98.50 99.94 80.25 95.68 97.00 99.53 93.75 98.93 92.38 98.52
ZeroScope 8 97.00 99.81 75.50 96.52 75.75 96.42 98.50 99.91 86.69 98.17
Text2Video-zero 16 99.50 99.99 60.70 95.13 65.55 93.55 63.27 95.38 72.25 96.01
Show1 16 97.86 99.69 95.93 99.04 93.73 98.13 94.30 98.88 95.45 98.93
ModelScope 16 100.00 100.00 81.25 95.41 98.36 99.55 95.36 99.19 93.74 98.53
ZeroScope 16 97.43 99.83 80.75 96.10 82.18 96.26 96.93 99.72 89.32 97.97
Text2Video-zero 24 99.50 100.00 57.57 95.14 55.7 94.10 60.18 96.32 68.23 96.39
Show1 24 97.80 99.98 94.36 99.10 89.66 97.49 93.73 98.93 93.88 98.87
ModelScope 24 100.00 100.00 74.68 95.57 97.00 99.32 90.73 99.30 90.60 98.54
ZeroScope 24 98.43 99.96 83.32 97.00 85.02 96.56 97.50 99.78 91.06 98.32
In our work, we selected 8 frames for generated video detection for two reasons: on the one hand, the current generation model still has this deficiency in generating long videos. Even the videos currently displayed by Sora can only generate about 300-600 frames, so we believe that 8 frames is an acceptable choice. On the other hand, in order to highlight the effectiveness of our method, we chose a challenging number of 8 frames, and a large number of experiments have also proven that our method indeed performs very well.
We conducted experiments on the impact of frames 8, 16, and 24 on detection, as shown in the table 6. We found that the performance of 16 frames is generally better than that of 8 frames, except for the models trained on the Text2Video zero subdataset, which is consistent with our previous evaluation results of several generation models. Text2Video zero has the worst time continuity. In addition, the effect will become worse at 24 frames. One possible reason is that as the sequence increases, the temporal artifacts of the generated video become more unique, which weakens the model’s generalizability.
Appendix F The Impact of Video Frame Sampling Interval.
In this section, we discussed the impact of sampling interval on detection results. In previous experiments, we usually took the first 8 frames of the video for detection.
Unlike previous experiments, this experiment was conducted on the latest generation model Sora, as Sora has made significant progress in generating frames compared to other generation models, making it easier for us to comprehensively discuss the impact of sampling strategies. As shown in the table 7, facing three different sampling strategies: 1) Take the first 8 frames; 2) Take 8 frames evenly from all frames; 3) Take 8 frames evenly from the first 300 frames; Our method did not show a significant decrease, which we attribute to two reasons: 1) The prompts we collected were comprehensive enough to cover the speed of video changes well. 2) The duration of the test video is still around 10 seconds, so the impact of sampling interval is not very significant. For longer generated videos, the performance degradation caused by changes in sampling intervals will be more severe.
Table 7: The impact of different sampling strategies
Training dataset Sampling strategy Sora
ACC AP
Text2Video-zero First 8 frames 96.87 99.70
Show1 First 8 frames 94.79 98.26
ModelScope First 8 frames 94.79 98.83
ZeroScope First 8 frames 89.58 97.92
Text2Video-zero Take 8 frames evenly from all frames 96.87 99.48
Show1 Take 8 frames evenly from all frames 94.79 98.61
ModelScope Take 8 frames evenly from all frames 94.79 98.96
ZeroScope Take 8 frames evenly from all frames 88.54 98.31
Text2Video-zero Take 8 frames evenly from first 300 frames 97.91 99.65
Show1 Take 8 frames evenly from first 300 frames 94.79 98.48
ModelScope Take 8 frames evenly from first 300 frames 93.75 98.74
ZeroScope Take 8 frames evenly from first 300 frames 90.62 98.22
Appendix G More Details on Sora.
G.1 Test data on Sora
Due to regional restrictions, we were not eligible to test Sora. But we tested all of the 48 videos officially displayed by Sora. The test dataset consists of 48 official videos released by Sora and 48 real videos from the original test set, the test on Veo are similar.
G.2 The detection performance of human testers on Sora
We invited three testers to evaluate the authenticity of 96 videos (48 real and 48 Sora generated) from the previously mentioned Sora test set. The experimental results are shown in the table 8.
Table 8: Compared to the test results of human testers
Training dataset Sora
TP+TN ACC
Tester 1 57(48+9) 59.37
Tester 2 66(47+19) 68.75
Tester 3 60(48+12) 62.5
Text2Video-zero 93 96.87
Show1 91 94.79
ModelScope 91 94.79
ZeroScope 86 89.58
The results of DeCoF in testing far exceed the level of human judgment. At the same time, we also need to point out that interviews with testers after testing revealed that their judgment on the videos generated by Sora was artificially high, as there were 4-5 flashy videos in the officially released videos that did not match the results that could appear in the real world, such as pirate ships in coffee cups. Meanwhile, according to the testers, their judgment is based on whether there is irregular movement (such as objects appearing out of thin air) or object threading (such as overlapping basketball and basket) and abnormal distortion (such as human fingers or limbs of moving animals) in the generated video.
Appendix H Spatial artifacts vs Temporal artifacts.
H.1 Detect generated videos from temporal artifacts
Although our method attempts to detect generated videos from temporal artifacts, there is a question: is it possible that CLIP: ViT has differences in the features between the generated video and the real video frames when extracting image features, resulting in better detection performance? In other words, is our method really detect generated videos from temporal artifacts?
To prove this, we repeated the previous probe experiment. As shown in the table 9, our model has also performed well in detecting test data without spatial artifacts. On the other hand, for the second set of test data in the previous probe experiment, our model has also successfully detected it. We believe that repeated stillness in multiple frames is also a type of temporal artifact to some extent.
Table 9: The proof of our method detect generated videos from temporal artifact
Method Video generation models
Temporal Spatial
ACC AP ACC AP
I3D 51.17 51.61 86.67 98.61
Slow 50.04 50.37 93.81 99.51
X3D 50.04 54.31 78.39 92.84
Mvit 50.48 55.29 72.16 92.59
Ours 98.46 99.80 98.46 99.66
H.2 Spatial artifacts vs Temporal artifacts.
In this paper, we attempt to detect generated video from the perspective of temporal artifacts, but this does not mean that it is impossible to detect generated video from spatial artifacts. In fact, the difficulty of detecting spatial artifacts in generated videos lies in how to extract more general artifacts, as different generation models generate videos with different resolutions, color brightness, contrast, and so on.
Appendix I Impact Statements
Our research focuses on detecting generated videos. We developed the DeCoF model and shared our dataset to support detecting generated videos. These works are crucial to protect digital content and prevent the spread of disinformation. Nevertheless, these tools have the potential to be misused, resulting in a rivalry between video generation and detection technology. We want to advocate for the ethical use of technology and foster research in the creation of tools for verifying media authenticity. We are confident that this will result in protecting the public from the harms of disinformation, improving the clarity and authenticity of information distribution, and ensuring the protection of personal privacy for individuals.
Appendix J Limitation
In this paper, we tackle an overlooked challenge, i.e., generated video detection, by proposing a universal detector to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models. However, there are various methods of video forgery, such as Deepfake and malicious editing, in which our methods may experience significant performance degradation or inapplicability. Taking Deepfake detection as an example, our method does not add additional detection modules to the face. In addition, Deepfake videos are often locally tampered with, and the generated videos are generated as a whole. Our work focuses on detecting generated videos, which has a more diverse range of forgery scenarios and more flexible forgery techniques.
Appendix K More Visual Examples